Evan is 6 years old. His sister is 15 years old.
How old will they both be when Evan is half as old as his sister?

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



(1) What is the difference in their ages?

(2) What will their ages be when Evan is half as old as his sister?

If you need a formal algebraic solution, re-post the question and say you want to see a formal algebraic solution.


Answer by ikleyn(52802)

.

            ALgebra solution


Let x = how many years to get the condition under the question.


Then at that time, Evan will be 6+x years old; His sister will be 15+x years old.


The condition is

    2*(6+x) = 15+x.


Simplify and solve

    12 + 2x = 15 + x

    2x - x = 15 - 12

     x     = 3.


The problem is just solved.


It will happen in 3 years.


Evan will be 6+3 = 9 years old;  his suster will be  15+3 = 18 years old.   ANSWER

Solved.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.